You are considering opening a new plant. The plant will cost

$ 99.9

million upfront. After​ that, it is expected to produce profits of

$ 31.6

million at the end of every year. The cash flows are expected to last forever. Calculate the NPV of this investment opportunity if your cost of capital is

6.3%.

Should you make the​ investment? Calculate the IRR. Use the IRR to determine the maximum deviation allowable in the cost of capital estimate to leave the decision unchanged.

Calculate the NPV of this investment opportunity if your cost of capital is

6.3%.

The NPV of this investment opportunity is

​$nothing

million. ​ (Round to one decimal​ place.)

Should you make the​ investment? ​ (Select the best choice​ below.)

A.

​Yes, because the project will generate cash flows forever.

B.

​No, because the NPV is less than zero.

C.

​Yes, because the NPV is positive.

D.

​No, because the NPV is not greater than the initial costs.

Calculate the IRR.

The IRR of the project is

nothing​%.

​ (Round to two decimal​ places.)

Use the IRR to determine the maximum deviation allowable in the cost of capital estimate to leave the decision unchanged.
Answer #1

All financials in $ million.

C0 = 99.9; C = 31.6; r = 6.3%

  1. NPV = - C0 + C / r = - 99.9 + 31.6 / 6.3% = 401.69
  2. IRR will be that r for which NPV will be zero i.e. - 99.9 + 31.6 / r = 0; Hence, IRR = r = 31.6 / 99.9 = 31.63%
  3. The maximum deviation allowable in the cost of capital estimate to leave the decision unchanged = IRR - r = 31.63% - 6.3% = 25.33% (i.e. an increase of 25.33%)
  4. Should you make the​ investment? ​ The correct answer is option C.​Yes, because the NPV is positive.